OpenStudy (anonymous):

solve the equation, x+9/7=10/7, for the given variable. write answer as a reduced fraction.

OpenStudy (texaschic101):

x + 9/7 = 10/7 -- subtract 9/7 from both sides x = 10/7 - 9/7 x = 1/7